Find the equation of the line through (6,-5) and parallel to 4x - 2y + 7 = 0
Answer by checkley75(3666)

4X-2Y+7=0
-2Y=-4X-7
Y=-4X/-2-7/-2
Y=2X+3.5 THIS LINE HAS A SLOPE OF 2
-5=6*2+b
-5=12+b
b=-5-12
b=-17 THUS THE EQUATION OF A PARALLEL LINE (SAME SLOPE) THROUGH (6,-5) IS
Y=2X-17